OpenStudy (anonymous):

solve the equation for n 8(n+7)-(1+3n)=4n

OpenStudy (petiteme):

Use distributive property first. So.. 8n+56-1-3n=4n Add like terms... 5n+55=4n Transpose 5n... 55=4n-5n 55=-n n=-55

OpenStudy (dan815):